The engine mechanism shown in Fig. 5 has crank OB = 50 mm and length of connecting rod


AB= 225 mm. The centre of gravity of the rod is at G which is 75 mm from B. The engine


speed is 200 r.p.m.


Fig. 5


For the position shown, in which OB is turned 45° from OA, Find 1. the velocity of G and the


angular velocity of AB, and 2. the acceleration of G and angular acceleration of AB.

1
Expert's answer
2021-12-24T10:19:02-0500

N=200rpm

OB=50mm=0.05m

AB=225mm=0.225m

GB=75mm=0.075m

"w_{OB}=\frac{2πN}{60}=\frac{2π*200}{60}=20.94rad/s"

"V_{OB}=OB*w_{OB}=0.05*20.94=1.047m\/s"

From velocity diagram,

"V_{A}=0.975m\/s"

"\frac{bg}{ba}=\frac{BG}{BA}"

"\frac{bg}{0.75}=\frac{75}{225}"

Hence, "bg=0.25m\/s"

Plot in velocity diagram to find VOG

"V_{OG}=0.975m\/s"